YoursTruly:

A bowling ball is moving down a bowling lane at 4 m/s. If the mass of the bowling ball is 7 kg, find its kinetic energy. Use the formula: KE = ½ mass x velocity^2 28 J 14 J 56 J 112 J

YoursTruly:

C?

YoursTruly:

@JustSaiyan

JustSaiyan:

56 J. KE = 1/2 of mass×velocity^2 so 7×4^2=112÷1/2=56

YoursTruly:

Thank you

JustSaiyan:

You're welcome.

JustSaiyan:

Here take a rusty star.